OpenStudy (anonymous):

Quick question...I'm given the matrix of the inverse, how do I directly calculate det(A) from this?

OpenStudy (anonymous):

det(inverse of A) = 1/det(A)

OpenStudy (anonymous):

seriously? cool, I'll try it out thanks!

OpenStudy (anonymous):

yes, you're welcome

OpenStudy (anonymous):

its a 4x4 matrix if thatchanges things

OpenStudy (anonymous):

it can be used on 4x4, 3x3, etc

OpenStudy (anonymous):

k cool

OpenStudy (anonymous):

but i dont think finding the determinant of 4x4 matrix easy. in my school the problem is usually 2x2 or 3x3

OpenStudy (anonymous):

Where does this equation come from? Is it unique, or does it derive from something else? It seems that it follows from A^-1=1/A..and the answer was right using the equation

OpenStudy (anonymous):

i forgot where it came from. it was on my notebook. my teacher gave it to me some months ago, along with : det(transpose of A) = det (A)

OpenStudy (anonymous):

yeah that one I knew, I'll have to look over my notes I think

OpenStudy (anonymous):

MathTy, I worked it out last night. The determinant of this 4x4 is 1! :D

OpenStudy (anonymous):

@MathTy: \[A*A^{-1}=I \implies \det(A*A^{-1})=\det(I) \implies \det(A)*\det(A^{-1})=1\]\[\implies \det(A^{-1})=1/\det(A).\] :)

OpenStudy (anonymous):

awesome I found it to be 1, but not nicely put like this. Does that mean that all invetible matrices have det=1?

OpenStudy (anonymous):

What do you mean "not nicely put like this"? xD That little proof just means that the product of the determinants of the regular matrix and the inverse matrix is always equal to 1. Not necessarily that the determinant of all invertible matrices is 1 -- this is just a clean exception. :P

OpenStudy (anonymous):

oh ok, by simple like this I just mean that to find the determinant I used cofactors to break it down into a3x3 matrix, then used rule of Sarrus to get the det of the 3x3, which came to 1